### v2

# ChromFound: Towards A Universal Foundation Model for Single-Cell Chromatin Accessibility Data

Abstract
The advent of single-cell Assay for Transposase-Accessible Chromatin using sequencing (scATAC-seq) offers an innovative perspective for deciphering regulatory mechanisms by assembling a vast repository of single-cell chromatin accessibility data. While foundation models have achieved significant success in single-cell transcriptomics, there is currently no foundation model for scATAC-seq that supports zero-shot high-quality cell identification and comprehensive multi-omics analysis simultaneously. Key challenges lie in the high dimensionality and sparsity of scATAC-seq data, as well as the lack of a standardized schema for representing open chromatin regions (OCRs). Here, we present ChromFound, a foundation model tailored for scATAC-seq. ChromFound utilizes a hybrid architecture and genome-aware tokenization to effectively capture genome-wide long contexts and regulatory signals from dynamic chromatin landscapes. Pretrained on 1.97 million cells from 30 tissues and 6 disease conditions, ChromFound demonstrates broad applicability across 6 diverse tasks. Notably, it achieves robust zero-shot performance in generating universal cell representations and exhibits excellent transferability in cell type annotation and cross-omics prediction. By uncovering enhancer-gene links undetected by existing computational methods, ChromFound offers a promising framework for understanding disease risk variants in the noncoding genome.
1 Introduction
The human genome harbors an extensive repository of open chromatin regions (OCRs) responsible for orchestrating precise spatial and temporal gene expression patterns [34]. Identifying OCRs at single-cell level is pivotal for understanding the regulatory landscape of individual cells, offering profound insights into cellular diversity and dynamic mechanisms of gene regulation [13, 9]. The single-cell Assay for Transposase-Accessible Chromatin using sequencing (scATAC-seq) [3] stands as the most prevalent genome-wide technique to enable the identification of OCRs at single-cell resolution. Its principal applications span cancer research, immunological studies, and neuroscience [53, 66, 64, 83]. Furthermore, with initiatives such as the Human Cell Atlas [56] and other significant studies [17, 87], the scale of scATAC-seq data is expanding rapidly, opening new possibilities and presenting novel challenges in the exploration of OCRs at the same time.
Pretrained language models (PLMs) have achieved great success in deciphering the complex language of single-cell RNA sequencing (scRNA-seq) data [67, 10, 28, 81], validating their capabilities fine-tuning for a variety of downstream tasks. Nevertheless, the development of generalizable foundation models for scATAC-seq remains under-explored, posing significant challenges for cellular representation learning and multi-task transfer. Here, we identify and discuss three major challenges.
(1). scATAC-seq data are inherently high-dimensional and sparse, with single cells spanning over a million OCRs but typically less than 1% showing accessibility. This sparsity complicates the utilization of Transformer [70] adopted in transcriptomics [67, 10, 28, 73, 19, 81]. Other methods such as transforming accessibility into gene activity [23, 62], filtering highly variable OCRs [23, 65, 21], or binarizing accessibility [29, 12, 85] always cause substantial information loss [63, 46, 5].
(2). scATAC-seq data are not formatted in a standardized feature space. Dynamic chromatin landscapes across diverse sources and platforms make dictionary-based tokenization for genes and nucleotides unsuitable for OCRs. All deep learning models for scATAC-seq [76, 75, 11, 85] require dataset-specific training, leading to poor zero-shot performance.
(3). current methods lack a unified integration of genomic information with chromatin accessibility profiles. Both genomic information defining regulatory loci and chromatin accessibility profiles revealing their activity are essential in scATAC-seq data. However, VAE-based models [76, 75, 11] process accessibility matrices independently, while models like scBasset, CellSpace, and SANGO [85, 65, 86] predict binary accessibility from genomic sequences. These methods struggle to generalize across diverse downstream tasks.
To address these challenges, we present ChromFound, a universal foundation model for single-cell chromatin accessibility data. Its hybrid architecture integrates a Mamba block for efficient long-range sequence processing with a self-attention block to capture local regulatory dependencies within ±200 kb of transcriptional start sites, aligning with the typical range of enhancer-promoter interactions identified by Hi-C analysis [1]. ChromFound incorporates biologically informed OCR tokenization encoding genomic coordinates and non-binary accessibility profiles. This component ensures scalability and alignment across diverse scATAC-seq datasets from varied tissues, platforms, and protocols. Pretrained on over 1.97 million single cells from more than 30 organs or tissues and 6 major disease categories (e.g., Alzheimer’s, Parkinson’s, leukemia, glioma), ChromFound leverages 1.86 trillion training tokens, surpassing the scale of existing models such as Geneformer[67], scGPT[10], and scFoundation[28], all of which use fewer than 1 trillion tokens.
To validate the effectiveness of ChromFound, we conduct comprehensive evaluations across multiple tasks and datasets. In cell representation, ChromFound outperforms baselines instead of additional training, achieving an average ARI improvement of 17.47% across 8 scATAC-seq datasets from 4 tissues. Additional results demonstrate its robustness to denoise technical noise in zero-shot settings. For cell type annotation, ChromFound significantly improves accuracy and macro F1 score across all evaluated datasets and tissues. In cross-omics prediction, it accurately infers gene expression from chromatin accessibility, surpassing all baselines across 5 datasets. Furthermore, in the K562 cell line, ChromFound effectively identifies gene-enhancer relationships and regulatory perturbation effects, highlighting its potential for interpretable cis-regulatory modeling.
Our main contributions can be summarized as follows:
(1) ChromFound, the first scATAC-seq foundation model, employs a genome-wide architecture along with a genome-aware tokenization and is pretrained on a large corpus with 1.97 million single cells.
(2) ChromFound provides a strong zero-shot cell representation from an epigenetic perspective, as well as showing consistent robustness to data sparsity and batch effect.
(3) ChromFound shows great transfer learning capabilities in cell annotation and cross-omics prediction, validating biological significance in inferring enhancer-gene links and perturbation responses.
2 Related Work
Single-cell foundation models for scRNA-seq data have seen significant advancements, supporting tasks such as gene function prediction, cell annotation, and drug response modeling. Geneformer [67], scGPT [10], and scFoundation [28] are trained by large-scale datasets, while scBERT [81] and CellPLM [73] focus on specific objectives based on pre-training language models such as imputation and cell communications. However, existing models are predominantly tailored for scRNA-seq data and fail to model scATAC-seq data effectively. GET [19] achieves a level of predictive precision comparable to experimental replicates by integrating chromatin accessibility data and genomic sequence information in a pseudo-bulk format instead of single-cell level.
Deep learning methods are widely used in scATAC-seq data analysis for cell clustering and cell annotation. SCALE [76] and its extension SCALEX [75] leverage the variational autoencoder [33] for latent feature extraction and multi-omics integration, while CASTLE [11] enhances cell type annotation through self-supervised learning. scBasset [85] connects the genomic sequence with binary accessibility and shows its strong ability to cell representation and batch integration. Cellcano  [45] and SANGO [86] employ a two-step MLP for cell type annotation. These models are impeded by their fundamental architectural limitations and the scarcity of available data, thereby restricting their scalability and generalizability across diverse datasets.
Figure 1: An Overview of ChromFound architecture. The methods of ChromFound are listed in Section 3. OCR tokenization are described in Section 3.1. The hybrid encoder layer is described in Section 3.2. Other pretraining details of datasets and implementation are included in Section 3.3. All evaluation results are detailed in Section 4.
3 Methods
3.1 OCR Tokenization
To encode chromatin accessibility data, we treat each OCR as a token by three essential components: chromosome embedding, positional embedding for genomic coordinates, and continuous accessibility embedding. The tokenization integrates these components to form a comprehensive representation of each OCR. In the following, we detail the encoding strategy for each component.
Chromosome Embedding: Chromosome identity is encoded using a learnable embedding lookup table. The size of the chromosome embedding lookup table is 25, with 24 corresponding to the number of human chromosomes and 1 designated for the padding token. We define the learnable embedding matrix $\mathbf{W}_{c}\in\mathbb{R}^{|\mathcal{C}|\times d_{\text{model}}}$ and the chromosome index $c_{j}$ of OCR $j$. ChromFound maps the $c_{j}$ to $\mathbf{E}_{c,j}$ by the following equation:
$\mathbf{E}_{c,j}=\mathbf{W}_{c}[c_{j},:].$ (1)
Positional Embedding for Genomic Coordinates: Since some studies [39, 79] have noted that state space models inherently handle sequential information through their recurrent nature, we use positional embedding to encode genomic coordinates instead of the sequence order. Specifically, the embedding of the starting position and the ending position for the OCR $j$ is computed as:
$\mathbf{E}^{s}_{p,j}=\begin{bmatrix}\sin\left(\frac{p^{s}_{j}}{\text{temp}\cdot 10000^{2k/d_{\text{model}}}}\right)\\ \cos\left(\frac{p^{s}_{j}}{\text{temp}\cdot 10000^{2k/d_{\text{model}}}}\right)\end{bmatrix},\quad\mathbf{E}^{e}_{p,j}=\begin{bmatrix}\sin\left(\frac{p^{e}_{j}}{\text{temp}\cdot 10000^{2k/d_{\text{model}}}}\right)\\ \cos\left(\frac{p^{e}_{j}}{\text{temp}\cdot 10000^{2k/d_{\text{model}}}}\right)\end{bmatrix}$ (2)
where $p^{s}_{j}$ and $p^{e}_{j}$ represents the starting and end position for the OCR $j$ on the GRCh38 reference genome, temp is a hyperparameter set to 100000 in ChromFound, and $d_{\text{model}}$ is the dimension of the embedding layer. The index $k$ spans $\{0,1,\dots,d_{\text{model}}/2\}$, with sine and cosine functions alternately assigned to even and odd dimensions of the embedding vector.
Accessibility Embedding: The accessibility value for the $j$th OCR, denoted as $X_{j}$, is a continuous scalar processed by the procedures detailed in Appendix C. To project the value $X_{j}$ into the hidden dimension, we use a linear transformation as $\mathbf{E}_{a,j}=\text{Linear}(X_{j})$.
OCR Representation: The unified representation $\mathbf{E}_{\text{OCR},j}\in\mathbb{R}^{1\times D}$ of the $j$th OCR is obtained by adding the previous embeddings together:
$\mathbf{E}_{\text{OCR},j}=\mathbf{E}_{c,j}+\mathbf{E}^{s}_{p,j}+\mathbf{E}^{e}_{p,j}+\mathbf{E}_{a,j}.$ (3)
3.2 Hybrid Encoder Layer
The ChromFound encoder layer transforms the OCR representation $\mathbf{E}_{\text{OCR}}\in\mathbb{R}^{L\times D}$ into a latent space, where $L$ denotes the sequence length and $D$ is the embedding dimension. As illustrated in Fig. 1, ChromFound incorporates a self-attention module (Section 3.2.1) to effectively capture local dependencies among OCRs, and integrates a Mamba layer (Section 3.2.2) to handle ultra-long OCR sequences. This architecture is motivated by biological insights, recognizing that both short- and long-range interactions among OCRs are critical to gene regulatory mechanisms [32, 52, 55].
3.2.1 Window Partition Self-Attention (WPSA)
Due to its computational complexity $O(L^{2})$, the vanilla attention mechanism [70] is not directly applicable to scATAC-seq data containing millions of OCRs. To address these challenges, ChromFound introduces an optimized self-attention layer named Window Partition Self-Attention (WPSA), applying attention mechanism to within a window divided by window size $W\ll L$.
Window Partition: Given the OCR representation $\mathbf{E}_{\text{OCR}}\in\mathbb{R}^{L\times D}$, $L$ may not be a multiple of $W$. We define the number of windows $N=\left\lceil\frac{L}{W}\right\rceil$ and pad $\mathbf{E}_{\text{OCR}}$ with zeros up to $N\times W$ along the dimension $L$. As a result, the padded sequence $\mathbf{E}_{\text{OCR}}^{\prime}\in\mathbb{R}^{(N\cdot W)\times D}$ is reshaped as $\mathbf{E}^{\prime\prime}_{\text{OCR}}\in\mathbb{R}^{N\times W\times D}$.
Self-Attention: In practice, we set the window size to $W=256$, which approximately spans 200 kb upstream and downstream of transcriptional start site (TSS). This setting is consistent with the scope of enhancer-promoter interactions suggested by high-resolution Hi-C analysis [1].
Window Restore: After applying self-attention within each window, we merge outputs of each window $\hat{\mathbf{E}}^{\prime\prime}_{\text{OCR}}$ back into $\hat{\mathbf{E}}_{\text{wpsa}}$:
$\hat{\mathbf{E}}_{\text{wpsa}}=\mathrm{concat}\bigl{[}\hat{\mathbf{E}}^{\prime\prime}_{\text{OCR},1},\ldots,\hat{\mathbf{E}}^{\prime\prime}_{\text{OCR},N}]\;\in\;\mathbb{R}^{(N\cdot W)\times D}.$ (4)
Furthermore, we remove any zero padding in $\hat{\mathbf{E}}_{\text{wpsa}}$ and get the WPSA output $\mathbf{E}_{\text{wpsa}}\in\mathbb{R}^{L\times D}$.
3.2.2 Mamba Block
Mamba [25] is capable of processing extended sequences due to the linear-time complexity of state space models (SSMs) [27, 26], making it particularly well suited for modeling high-resolution scATAC-seq data. When combined with the WPSA module, which amplifies local signals within predefined windows, the Mamba block enables effective integration of both short- and long-range dependencies among OCRs across the genome. This hybrid architecture mitigates the degradation of long-range information typically caused by relying solely on zero-order discretization [51, 84]. To reduce parameter count and refine representations, we project $\mathbf{E}_{\text{wpsa}}$ to a lower-dimensional space $\mathbb{R}^{L\times D_{\mathrm{low}}}$ and then back, which is define as:
$\mathbf{E}_{\mathrm{down}}=\mathbf{E}_{\text{wpsa}}W_{\mathrm{down}},\quad\mathbf{E}_{\mathrm{mamba}}=\mathrm{Mamba}(\mathbf{E}_{\mathrm{down}}),\quad\mathbf{E}_{\mathrm{up}}=\mathbf{E}_{\mathrm{mamba}}W_{\mathrm{up}},$ (5)
where $W_{\mathrm{down}}\in\mathbb{R}^{D\times D_{\mathrm{low}}}$ ($D_{\mathrm{low}}<D$) and $W_{\mathrm{up}}\in\mathbb{R}^{D_{\mathrm{low}}\times D}$.
3.2.3 Overall Workflow
The overall workflow of the encoder layer is illustrated in Fig. 1 and proceeded as:
$\mathbf{E}_{\text{out}}=\mathbf{E}_{\text{OCR}}+W_{\mathrm{up}}\cdot\text{Mamba}(W_{\mathrm{down}}\cdot\text{WPSA}(\text{RMSNorm}(\mathbf{E}_{\text{OCR}})))$ (6)
3.3 Pre-training
Datasets: We assemble a large-scale scATAC-seq dataset comprising over 2.64 million cells detailed in Table 5. For model pretraining, we select a representative subset of 1.97 million cells from more than 30 distinct organs and tissues. To rigorously evaluate the generalization ability of ChromFound, we additionally compile a benchmark dataset containing 0.67 million cells from diverse sources, including tissues such as bone and retina excluded from pretraining.
Implementation: ChromFound is trained for 5 epochs over 80 hours on a compute cluster comprising 4 machines with a total of 32 NVIDIA A100 GPUs. The effective batch size is set to 128. We use the AdamW optimizer [42] with a maximum learning rate of $5\times 10^{-5}$. The embedding dimension and the hidden size of WPSA $D$ are set to 128, while dimension $D_{\text{low}}$ of the Mamba block is set to 32. The model architecture consists of 4 stacked encoder layers in total.
Training objective: ChromFound adopts a masking strategy that simultaneously targets both zero and non-zero OCR values. Inspired by xTrimoGene [22], the setting encourages the model to predict non-binary chromatin accessibility across the entire profile, mitigating the risk of learning representations dominated by zero entries. The MLP layer is employed as the decoder to reconstruct the original input $X_{i}$. The reconstruction output of cell $i$ is denoted as $\hat{X}_{i}$. The pretraining objective is the mean squared error (MSE) computed over the masked positions $\mathcal{M}_{i}$ in each cell $i$:
$\mathcal{L}_{mse,i}=\frac{1}{|\mathcal{M}_{i}|}\sum_{i\in\mathcal{M}_{i}}\left(X_{i}-\hat{X}_{i}\right)^{2}.$ (7)
Table 1: Results of cell clustering tasks. Result style: best, second best, relative gains.
Dataset Tissue Model ARI($\uparrow$) FMI($\uparrow$) NMI($\uparrow$) AMI($\uparrow$)
Morabito130K[49] (batch 1) Cortex SCALE[76] 0.5460±0.0074 0.6896±0.0042 0.7282±0.0016 0.7275±0.0016
SCALEX[75] 0.4930±0.0022 0.6505±0.0013 0.6958±0.0001 0.6949±0.0001
CASTLE[11] 0.4425±0.0029 0.6109±0.0019 0.7086±0.0012 0.7077±0.0012
scBasset[85] 0.5039±0.0039 0.6747±0.0022 0.6674±0.0005 0.6647±0.0005
ChromFound 0.6890±0.0387 0.7943±0.0144 0.7779±0.0037 0.7760±0.0038
Gains(%) 26.20 15.17 6.81 6.66
Morabito130K[49] (batch 2) Cortex SCALE[76] 0.5306±0.0048 0.6394±0.0032 0.6221±0.0007 0.6200±0.0007
SCALEX[75] 0.3905±0.0041 0.5252±0.0027 0.5361±0.0028 0.5336±0.0029
CASTLE[11] 0.3743±0.0016 0.5087±0.0010 0.4636±0.0021 0.4608±0.0022
scBasset[85] 0.5373±0.0024 0.6424±0.0015 0.6659±0.0003 0.6641±0.0003
ChromFound 0.6278±0.0043 0.7156±0.0025 0.7321±0.0016 0.7299±0.0016
Gains(%) 16.83 11.40 9.91 9.95
Kuppe139K[36] (donar av3) Heart SCALE[76] 0.4475±0.0015 0.5654±0.0011 0.6005±0.0003 0.5978±0.0003
SCALEX[75] 0.3819±0.0007 0.5061±0.0005 0.4829±0.0000 0.4797±0.0000
CASTLE[11] 0.3886±0.0017 0.5118±0.0110 0.4765±0.0001 0.4729±0.0001
scBasset[85] 0.4553±0.0022 0.5729±0.0016 0.6539±0.0006 0.6516±0.0006
ChromFound 0.5828±0.0052 0.6757±0.0034 0.7207±0.0008 0.7187±0.0008
Gains(%) 28.01 17.94 10.21 10.31
Kuppe139K[36] (donar av10) Heart SCALE[76] 0.4517±0.0087 0.6596±0.0040 0.5397±0.0041 0.5391±0.0041
SCALEX[75] 0.4742±0.0105 0.6770±0.0055 0.5040±0.0079 0.5033±0.0080
CASTLE[11] 0.4739±0.0000 0.6786±0.0000 0.5204±0.0000 0.5197±0.0000
scBasset[85] 0.5823±0.0111 0.7459±0.0050 0.6780±0.0001 0.6776±0.0001
ChromFound 0.6774±0.0445 0.8109±0.0169 0.7369±0.0137 0.7365±0.0138
Gains(%) 16.34 8.71 8.68 8.70
Liang154K[38] (sample D026_13) Retina SCALE[76] 0.5512±0.0386 0.7285±0.0075 0.7172±0.0051 0.7168±0.0051
SCALEX[75] 0.5069±0.0048 0.7241±0.0022 0.6759±0.0005 0.6755±0.0005
CASTLE[11] 0.2424±0.0032 0.5136±0.0022 0.5298±0.0033 0.5292±0.0033
scBasset[85] 0.4330±0.0049 0.6699±0.0029 0.6903±0.0015 0.6898±0.0015
ChromFound 0.6668±0.0500 0.8149±0.0169 0.7644±0.0093 0.7641±0.0093
Gains(%) 20.98 11.86 6.59 6.60
Liang154K[38] (sample D19D008) Retina SCALE[76] 0.6129±0.0131 0.7465±0.0113 0.7981±0.0020 0.7993±0.0058
SCALEX[75] 0.5424±0.0041 0.6872±0.0023 0.6976±0.0019 0.6971±0.0019
CASTLE[11] 0.4272±0.0052 0.5982±0.0034 0.6614±0.0034 0.6608±0.0034
scBasset[85] 0.6027±0.0139 0.7313±0.0066 0.7881±0.0010 0.7878±0.0010
ChromFound 0.6688±0.0264 0.7767±0.0123 0.8183±0.0032 0.8179±0.0032
Gains(%) 9.12 4.05 2.53 2.34
PBMC169K[15] (batch VIB_10xv1_1) PBMC SCALE[76] 0.6158±0.0048 0.6959±0.0031 0.7632±0.0010 0.7623±0.0010
SCALEX[75] 0.6121±0.0048 0.6927±0.0015 0.7593±0.0008 0.7584±0.0008
CASTLE[11] 0.5530±0.0018 0.6439±0.0012 0.7263±0.0004 0.7252±0.0004
scBasset[85] 0.6240±0.0001 0.7019±0.0001 0.7660±0.0000 0.7651±0.0000
ChromFound 0.6953±0.0035 0.7601±0.0023 0.7860±0.0007 0.7852±0.0008
Gains(%) 11.44 8.29 2.61 2.62
PBMC169K[15] (batch BIO_ddseq_1) PBMC SCALE[76] 0.4359±0.0181 0.6250±0.0101 0.5594±0.0045 0.5587±0.0045
SCALEX[75] 0.4087±0.0197 0.6063±0.0106 0.5519±0.0066 0.5512±0.0066
CASTLE[11] 0.3504±0.0042 0.5555±0.0026 0.5512±0.0013 0.5506±0.0013
scBasset[85] 0.4362±0.0012 0.6223±0.0008 0.5577±0.0008 0.5571±0.0008
ChromFound 0.4835±0.0082 0.6604±0.0042 0.5950±0.0036 0.5944±0.0036
Gains(%) 10.84 5.66 6.37 6.39
4 Experiments
To systematically evaluate the performance of ChromFound, we conduct experiments on various tissues and datasets. We first assess its zero-shot representation capabilities, demonstrating strong robustness to both data sparsity and batch effects. We then evaluate the transferability in downstream tasks, including cell type annotation and cross-omics prediction. Finally, we validate the biological utility of ChromFound by accurately inferring gene-enhancer regulatory links and gene expression responses to enhancer perturbations. All scATAC-seq datasets used in experiments are preprocessed according to the procedures detailed in Appendix C.
4.1 Cell representation
ChromFound generates low-dimensional cell representations by applying PCA to the encoder outputs after the pooling layer. Details of the implementations are provided in Appendix A.1. For comparison, we include scBasset [85] and three VAE-based models [76, 75, 11], which require self-supervised learning on each specific dataset. We conduct experiments on eight datasets from four tissue types and evaluate performance using four clustering metrics to ensure fair and reliable comparison.
All results presented in Table 1 reveal that ChromFound outperforms existing SOTA methods in all tissues and metrics, with average improvements of 17.47%, 10.39%, 6.72% and 6.69% in ARI, FMI, NMI and AMI, respectively. The greater improvements observed in ARI compared to NMI and AMI infer the robustness of ChromFound to technical noise such as low read depth and batch effects, as ARI is particularly sensitive to local structures within the cell representation space. We discuss ChromFound’s ability to denoising low count and batch effect as below.
Denoising low count: Some studies [5, 37, 65] have acknowledged that the prevalence of missing signals in scATAC-seq makes standard analyses challenging. Although all datasets in Table 1 already exhibit high sparsity (~99%), we further simulate increasingly sparse conditions by downsampling accessible OCRs to 10%, 20%, 30%, 40% and 50% of their original counts. We adopt the same benchmarking methods and clustering metrics as listed in Table 1. As shown in Fig. 2, ChromFound consistently maintains stable performance across all levels of downsampling, with its advantage over baseline methods becoming increasingly pronounced as the sparsity intensifies.
Denoising batch effect: Denoising batch effects remains a major challenge in scATAC-seq analysis [47, 65, 44]. We evaluate ChromFound on datasets from 4 tissues using biological conservation metrics [44] consistent with scGPT [10]. In addition to SCALEX [75] and scBasset [85], we include five specialized batch integration methods [41, 31, 35, 77, 40] for comparison. As shown in Table 2, ChromFound in zero-shot settings outperforms existing methods by more effectively mitigating batch effects while better preserving biologically significant variation.
Figure 2: Results of denoising low count. The left y-axis shows the absolute metric values (scatter plot) and the right y-axis indicates the relative gains (%) over the second-best methods (bar plot).
Table 2: Results of denoising batch effect. Result style: best, second best, relative gains. AvgBIO is the average of $\text{ARI}_{\text{cell}}$, $\text{NMI}_{\text{cell}}$ and $\text{ASW}_{\text{cell}}$ measuring biological consistency. AvgBATCH is computed as the average of $\text{ASW}_{\text{batch}}$ and GraphConn to summarize the batch mixing performance.
Model
Bone To326K[68]
batch 43/44
Heart Kuppe139K[36]
donar av3/av10
PBMC 169K[15]
HAR ddseq 1/2
Cortex Morabito130K[49]
batch 1/2
AvgBIO($\uparrow$) AvgBATCH($\uparrow$) AvgBIO($\uparrow$) AvgBATCH($\uparrow$) AvgBIO($\uparrow$) AvgBATCH($\uparrow$) AvgBIO($\uparrow$) AvgBATCH($\uparrow$)
scVI[41] 0.3713 0.9026 0.7303 0.8448 0.6101 0.7932 0.7238 0.9217
Scanorama[31] 0.4386 0.8722 0.5320 0.7779 0.6036 0.7964 0.6895 0.9243
Harmony[35] 0.3789 0.8558 0.5119 0.8390 0.6232 0.7923 0.7152 0.9101
scANVI[77] 0.4351 0.8479 0.6810 0.8401 0.6028 0.8035 0.7185 0.9208
Liger[40] 0.1992 0.6471 0.7286 0.8019 0.6051 0.8009 0.6822 0.9208
SCALEX[75] 0.1778 0.9177 0.7596 0.8415 0.5989 0.8004 0.4789 0.8728
scBasset[85] 0.3650 0.9208 0.6216 0.8591 0.6024 0.8014 0.7045 0.9301
ChromFound 0.6408 0.9289 0.8180 0.8679 0.6443 0.8217 0.7440 0.9565
Gain(%) 46.07 0.88 7.69 1.03 3.37 2.25 2.78 2.84
4.2 Cell Type Annotation
Computational cell type identification is a fundamental task in single-cell omics analysis. To adapt ChromFound for this task, we introduce a feature projection head followed by MLP layers as described in Appendix F.1. Our methods are compared to three strong baseline models designed for cell type annotation of scATAC-seq, Cellcano [45], EpiAnno [6] and SANGO [86]. The variant of ChromFound trained from scratch is also included to assess the contribution of pretraining. Fig. 3 highlights that ChromFound consistently outperforms all benchmark methods, achieving the highest performance in macro F1 score and demonstrating its ability to accurately annotate diverse cell types. The confusion matrices on PBMC169K [15] for ChromFound and CellCano [45] are shown in Fig. 5. The results indicate that Cellcano struggles to correctly classify rare cell types such as natural killer cells, CD16+ monocytes, and dendritic cells, likely due to their low abundance. In contrast, ChromFound significantly improves the classification accuracy for these cell types, enabling a more precise understanding of immune cell composition and dynamics.
Figure 3: Results of cell type annotation, evaluated by Accuracy (left) and macro F1 Score (right).
4.3 Cross-omics Prediction
Although recent advances have enabled the simultaneous profiling of multiple assays, most single-cell datasets remain isolated, posing significant challenges for effective multi-omics analysis. To explore the potential of ChromFound in multi-omics prediction, we utilize scATAC-seq data as the source modality and scRNA-seq data as the target modality. Implementation details are provided in Appendix A.3. For comprehensive evaluation, we benchmark on five paired scATAC-seq and scRNA-seq datasets [88, 68, 43, 4, 89] using two standard correlation metrics. Three representative baseline methods [74, 82, 72] are included with default configurations from the DANCE toolkit [16]. As shown in Table 3, ChromFound consistently outperforms all baselines across datasets in both PCC and CCC, highlighting its strong capability in cross-omics prediction. These results highlight the potential of ChromFound to facilitate biological analyses such as inferring gene-enhancer relationships and predicting transcriptional response to enhancer perturbation. The two biological applications are discussed in detail below.
Table 3: Cross-omics prediction results. Result style: best, second best, relative gains. ChromFound* stands for the version of ChromFound without loading the pretrained checkpoint.
Model Cortex Zhu45K[88] Bone To326K[68] BMMC multiome 2021[43] BMMC atac2gex 2022[4] Cell lines Zhu11K[89]
[…]
BABEL[74] 0.7975 0.7687 0.8020 0.7223 0.3854 0.3695 0.8901 0.8329 0.9196 0.8608
CMAE[82] 0.7973 0.7525 0.7986 0.6683 0.3435 0.3204 0.8983 0.7990 0.9136 0.8539
scMM[48] 0.7793 0.7325 0.7586 0.6513 0.3718 0.2983 0.8321 0.7027 0.8587 0.8062
scMoGNN[72] 0.8001 0.7210 0.7957 0.7233 0.4168 0.3911 0.9124 0.8655 0.7976 0.7263
ChromFound* 0.7864 0.7674 0.7887 0.6919 0.4213 0.3989 0.9279 0.8704 0.8992 0.8501
ChromFound 0.8316 0.8064 0.8304 0.7472 0.4249 0.4032 0.9293 0.8818 0.9449 0.9071
Gain(%) 4.10 4.67 3.54 3.30 1.94 3.09 1.85 1.88 2.67 5.11
Biological Application: Predicting enhancer-gene link and perturbation response
Enhancer elements in the human genome harbor thousands of genetic variants that regulate gene expression and contribute to the risk of common diseases. In our approach detailed in Appendix A.4, we simulate OCR knockdown and estimate the resulting changes in gene expression. The absolute change reflects the likelihood of an enhancer-gene regulatory link, while the relative change captures the direction and strength of the transcriptional response. We perform experiments on 141 K562 cells from the test split in Zhu11K [89], using the cross-omics model trained as evaluated in Table 3. Fulco4K [89] provides the ground truth of enhancer-gene links and quantitative effects through CRISPRi perturbations in K562 human erythroleukemia cells. For clarity, the evaluation focuses on the COPZ1 and HNRNPA1 genes, which are known to be involved in cancer-related processes such as stress adaptation and RNA splicing regulation [61, 7]. BABEL [74] and CMAE [82] are included as benchmark methods due to their strong performance in cross-omics prediction in Zhu11K [89]. As shown in Fig. 4, ChromFound achieves the highest AUC scores with 0.77 for COPZ1 and 0.61 for HNRNPA1, accurately identifying enhancer-gene links. Fig. 4 further shows that ChromFound better predicts transcriptional responses to enhancer perturbations. In contrast, BABEL [74] and CMAE [82] filter for the top 10,000 highly variable OCRs [16], leading to zero predictions of most enhancer perturbations. This limitation is evident in both Fig.4 and Fig.4, highlighting the advantage of ChromFound’s genome-wide modeling to support a more comprehensive understanding of regulatory mechanisms.
Figure 4: (a) ROC curves illustrate the performance of enhancer-gene link prediction, where true positive rates are plotted against false positive rates across varying prediction thresholds. Ground truth labels are defined based on the significance (FDR < 0.05) of expression changes after enhancer perturbation. (b) Scatter plots represent a specific enhancer perturbation. The y-axis shows the predicted scores of average gene responses of enhancer perturbation, while the x-axis represents the real effects of post-perturbation. Both magnitudes are rescaled to the unit norm.
4.4 Ablation study
Table 4: Ablation study. Red row indicates the full model of ChromFound. The other colored rows correspond to the topics discussed in Section 4.4.
Genome
info
WPSA Mamba
Data size
# of OCRs
per cell
FLOPs(10e9)
per cell
ARI($\uparrow$) FMI($\uparrow$) NMI($\uparrow$) AMI($\uparrow$)
1.97 440000 19.72 0.6953±0.0035 0.7601±0.0023 0.7860±0.0007 0.7852±0.0008
1.97 440000 19.72 0.6452±0.0149 0.7095±0.0098 0.7227±0.0021 0.7301±0.0017
1.97 440000 19.72 0.5897±0.0086 0.6401±0.0113 0.6419±0.0012 0.6476±0.0015
1.97 440000 19.72 0.3075±0.0019 0.3098±0.0011 0.3101±0.0011 0.3326±0.0013
0.2 440000 19.72 0.6539±0.0057 0.7261±0.0037 0.7593±0.0012 0.7583±0.0012
0.02 440000 19.72 0.6142±0.0053 0.6995±0.0034 0.7354±0.0015 0.7345±0.0015
1.97 220000 9.87 0.4012±0.0047 0.4500±0.0031 0.4535±0.0007 0.4425±0.0007
1.97 110000 4.93 0.3276±0.0051 0.3589±0.0048 0.3674±0.0002 0.3668±0.0002
In this section, we conduct an ablation study to address three questions using the PBMC169K [15], sampled from batch VIB_10xv1_1. The results of cell representation are summarized in Table 4.
Question 1: Does our proposed ChromFound architecture contribute to performance?
We evaluate the individual impact of each key module in our model, including genome-aware tokenization (Section 3.1), the WPSA layer (Section 3.2.1) and the Mamba block (Section 3.2.2). In rows 2-4 of Table 4, we observe a clear performance drop in all clustering metrics compared to row 1, demonstrating that each component contributes complementarily to overall performance.
Question 2: Does scaling to large-scale data provide performance gains?
As shown in rows 5-6 of Table 4, reducing the data size from 1.97 million to 0.2 million and 20 thousand cells leads to consistent degradation in all evaluation metrics, confirming that large-scale data learning facilitates better generalization.
Question 3: Is long-context modeling necessary for high-quality representations?
As seen in rows 7-8 of Table 4, reducing the input length by half and a quarter significantly hurts performance, highlighting the importance of long-context modeling for dynamic OCR landscapes.
5 Conclusion
In this work, we present ChromFound, to the best of our knowledge, the first foundation model specifically for scATAC-seq data. To address the inherent challenges of high-dimensionality, sparsity, and dynamic chromatin landscapes, we utilize a hybrid architecture that integrates short- and long-range dependencies with a genome-aware tokenization. Trained on 1.97 million single-cell profiles comprising over 30 human tissue types and 6 disease categories, ChromFound achieves SOTA performance in 6 downstream tasks, showing great robustness in zero-shot cell representation. Its biological significance is further demonstrated by its ability to accurately infer transcriptional responses to enhancer perturbations. In the future, we aim to explore the broader applicability of ChromFound in the large-scale mapping of enhancer-gene regulatory interactions, a critical yet largely uncharted component of the cis-regulatory landscape of the human genome.
Appendix A Experiments Details
A.1 Cell Clustering
Since ChromFound is configured in a zero-shot mode for the cell clustering task, no additional fine-tuning architecture or training parameters need to be specified. We directly propagate the data through the encoder structure of ChromFound, extracting the encoder output as the final representation. After performing a pooling layer on the hidden dimension, we apply Principal Component Analysis (PCA) to reduce the dimensionality of the output to 50. Based on this, we compute the metrics for cell clustering and generate the UMAP clustering visualization. The benchmark methods are implemented with the default parameters provided by their source codes.
We compute the inference speed of ChromFound on a node equipped with 32 CPU cores, 256 GB of memory, and one NVIDIA A100 GPU. The input scATAC-seq dataset with 232,354 OCRs. The maximum CPU memory usage is 23.9 GB and the maximum GPU memory consumption is 12.7 GB out of 80 GB available. The inference speed with batch size 4 is 0.97 seconds per batch.
A.2 Cell Type Annotation
Building upon our four-layer pretrained model, we extract the encoder output as a general-purpose representation for downstream cell type annotation. To align the model with this task, we incorporate additional fine-tuning layers. Specifically, the hidden dimension produced by ChromFound’s encoder is expanded from 128 to 256 before being projected down to a single dimension, resulting in a pooling tensor. Subsequently, additional multilayer perceptron (MLP) layers are employed to predict the logits corresponding to cell types. We freeze the decoder parameters from pretraining while enabling gradient flow through the rest of the pretrained backbone so that it can adapt to the new classification objective. Dropout mitigates overfitting by randomly zeroing hidden units, while LayerNorm helps stabilize and accelerate fine-tuning. The benchmark methods are implemented with the default parameters provided by their source codes.
For experimental settings, we divide the training data into 90% for training and 10% for validation. We train for 20 epochs using the AdamW optimizer with an initial learning rate of $5\times 10^{-4}$ and a linear warm-up schedule of 50 steps. The best model is chosen based on validation set performance and evaluated on the test set to obtain the reported accuracy and macro F1 score. All experiments are conducted on a single machine equipped with four NVIDIA A100 GPUs.
A.3 Cross-omics Prediction
We employ a two-layer Multilayer Perceptron (MLP) to perform the modality prediction task, with the hidden layer size set to 1024. The learning rate is configured at 1e-4, and AdamW is utilized as the optimizer. A total of 10 epochs are conducted for training. All benchmark methods are implemented with the default parameters in the DANCE package [16]. All experiments are conducted on a single machine equipped with four NVIDIA A100 GPUs. We divide the whole dataset into 80% for training, 10% for evaluation and 10% for testing.
A.4 Biological Application
To validate the biological relevance of predicted enhancer-gene interactions, we simulate enhancer knockdown using our cross-omics model trained on the Zhu11K dataset [89]. Specifically, we select 141 K562 cells from the test split and iteratively set the accessibility of each candidate enhancer to zero, mimicking CRISPRi-mediated repression. The model then infers post-perturbation gene expression, and the change relative to the unperturbed state is computed and averaged across all cells. The absolute magnitude of this change indicates the likelihood of a regulatory interaction, while the signed change reflects its direction and strength.
We focus on two cancer-related genes, COPZ1 and HNRNPA1, which are extensively characterized in the Fulco4K CRISPRi dataset [89]. Each gene is associated with 117 candidate enhancers, among which 6 (COPZ1) and 10 (HNRNPA1) are experimentally validated. To resolve discrepancies in genomic coordinates between Fulco4K [20] and Zhu11K [89], we perform enhancer mapping using bedtools [54]. In total, the simulation spans 230,819 enhancers and 17,476 genes.
To evaluate performance, we compute the area under the ROC curve seen in Table 4 between the absolute expression change and the binary labels of enhancer-gene links tested with a statistical significance on gene expression at a false discovery rate (FDR) < 0.05 [20]. ChromFound achieves the highest accuracy, with AUCs of 0.77 for COPZ1 and 0.61 for HNRNPA1, demonstrating its ability to identify functional regulatory links. Additionally, we calculate the Pearson correlation between signed responses and CRISPRi-measured quantitative effects plotted in Table 4.
The Benchmark methods BABEL [74] and CMAE [82] are included for comparison due to their strong cross-omics performance on Zhu11K. However, both rely on filtering the input scATAC-seq profiles to the top 10,000 highly variable OCRs [16], which excludes the majority of candidate enhancers from Fulco4K. As a result, these methods often fail to produce meaningful predictions for enhancer perturbations, yielding near-zero expression changes in most cases. This limitation is evident in both the ROC and scatter plots, underscoring the advantage of ChromFound’s genome-wide modeling in capturing fine-grained regulatory effects at scale.
Appendix B Details of Datasets
We collect a large-scale human scATAC-seq dataset that includes more than 2.65 million cells and 1.75 trillion tokens. ChromFound for pretraining is based on two fundamental datasets, the human atlas [87] and the fetal atlas [17], which contribute 1.32 million cells spanning multiple human organs and serve as the primary sources for pretraining ChromFound. We accept the CRCh38 as reference genome across all datasets to maintain uniformity, converting datasets originally in hg19 to hg38 when necessary. The resources of all datasets are detailed in Table 5.
Appendix C Pre-processing of scATAC-seq Data
We implement a pre-process pipeline to scATAC-seq data, addressing sparsity and dimensionality to ensure high-quality cells and OCRs for model pretraining and downstream tasks. Pre-processing is performed after the train/test splits in pretraining and downstream tasks, preventing any information leakage. All datasets are saved in .h5ad format with the OCR coordinates of genome in var schema. We detail the pre-processing methods in the following.
Cell filtering: We filter cells with non-zero counts lower than 1000 or higher than 60000 for every dataset.
OCR filtering: We compute the non-zero counts of each OCR by cell types and filter OCRs with non-zero counts lower than 5% in all cell types, ensuring that long-tail cell types remain heterogeneous.
Normalization and Log Transformation: To stabilize variance and scale the data, we apply total count normalization followed by log transformation.
Appendix D More related works about Mamba and hybrid architecture
Mamba-based models, such as Mamba [25] and Mamba-2 [14], replace self-attention with selective State Space Models (SSMs) [60, 27, 26], achieving linear complexity while maintaining strong performance. Recently, Mamba has been applied in biology and genomics such as RNA foundation models [18] and DNA sequence modeling [59], efficiently handling sparse and long-range dependencies in biological data.
We have noticed that some works have started to use hybrid transformer-mamba structures for modeling time series and image data. After conducting some theoretical research and analysis, we found that the existing hybrid structures are not suitable for scATAC-seq modeling. Therefore, ChromFound has been meticulously designed for scATAC-seq data, with fundamental differences from existing hybrid structures. Below, we highlight two important papers for a detailed theoretical comparison.
1. The SST [80] uses transformers and mamba to separately model short-range and long-range dependencies, then concatenates the features of short-range and long-range interactions. This resembles a parallel stacking structure. In contrast, ChromFound employs a sequential stacking approach, where data first passes through the transformer for short-range feature interactions and then moves to mamba for long-range learning.
2. MambaVision [30] integrates Vision Transformers with Mamba and is tailored for vision applications. ChromFound, however, has been designed specifically to accommodate the local strong interactions present in scATAC-seq data. We introduced a novel window partition self-attention module to better capture these interactions.
Table 5: Dataset Information
Dataset Tissue Disease Cell Numbers OCR Numbers Genome GSE/GSM ID HumanAtlas [87] Atlas Health 615,998 1,154,611 hg38 GSE184462 FetalAtlas [17] Atlas Health 707,043 1,154,646 hg38 GSE149683 PBMCBenchmark [15] Peripheral Blood Health 169,047 412,490 hg38 GSE194028 Cortex130K [49] Cortex Alzheimer’s disease 130,419 219,070 hg38 GSE174367 Cortex2K [50] Cortex Health 2,174 292,156 hg38 GSE174226 Zhu45K [88] Cortex Health 45,549 304,034 hg38 GSE204684 PBMC9K [24] Peripheral Blood Leukemia 9,215 108,344 hg38 GSE139369 BMMC11K [24] Bone Marrow Health 11,384 452,004 hg19 GSE194122 Rubin [57] Epidermis Health 288 94,633 hg19 GSE116428 Xu12K [78] Breast Breast cancer 7,202 87,851 hg38 GSM4798906 5,642 76,850 GSM4798907 Wang23K [71] Brain Glioma 6,284 79,730 hg38 GSM4119513 129 1,169 GSM4119514 5,213 210,434 GSM4119515 5,519 183,847 GSM4119516 2,229 40,907 GSM4119517 3,628 93,121 GSM4119518 Buenrostro3K [2] Bone Marrow Health 2,953 491,437 hg19 GSE96769 Corces0.5K [8] Bone Marrow/ Peripheral Blood Leukemia 576 590,650 hg19 GSE74310 Corces70K [9] Brain Alzheimer’s and Parkinson’s diseases 70,631 444,747 hg38 GSE147672 Satpathy110K [58] Bone Marrow Health 63,882 571,400 hg19 GSE129785 Peripheral Blood Health 4,146 238,616 Peripheral Blood Health 4,786 152,367 TME Cancer 37,818 580,789 Ziffra77K [90] Forebrain Health 77,354 459,953 hg38 GSE163018 Kuppe139K [36] Heart Myocardial infarction 139,835 429,828 hg38 Liang154K [38] Retina Health 154,775 264,833 hg19 GSE226108 To326K [68] Bone Health 326,532 530,751 hg38 Zhu11K [89] Cell Line N/A 11,632 258,044 hg19 GSE118912 10x PBMC Datasets Peripheral Blood Health 10,247 90,686 hg19 10k v1.11 9,688 144,023 10k v2.02 4,623 135,377 5k v2.03 1,004 82,579 1k v2.04 484 65,908 500 v2.05
Appendix E Discussion of comparing OCR tokenization with other peak-calling algorithms
Below, we discuss about its experimental justification and practical utility compared to approaches relying on consistent peak-calling pipelines.
1. Experimental validation and ablation study demonstrate the efficacy of genome-aware OCR tokenization: Our genome-aware tokenization framework integrates three biologically grounded components: chromosome identity, genomic coordinates, and accessibility values. By explicitly encoding positional and genomic context, ChromFound adapts to OCRs derived from heterogeneous peak-calling pipelines without requiring a unified feature space. The ablation study in 4 empirically validates this design: the inclusion of genome-aware tokenization improves clustering performance compared to a baseline lacking genomic embeddings.
2. Genome-aware tokenization reduces preprocessing complexity: A critical advantage of ChromFound lies in its ability to operate directly on precomputed count matrices (e.g., .h5ad files), eliminating dependencies on raw fragment files or reprocessing through specific peak-calling pipelines. Publicly available scATAC-seq datasets often only provide processed matrices, rendering methods reliant on unified peak sets impractical. ChromFound circumvents these limitations, ensuring immediate applicability to datasets lacking raw sequencing data. Even when fragment files are accessible, reprocessing data through unified pipelines (e.g., ENCODE) imposes prohibitive burdens. Such workflows require extensive data preprocessing, and substantial computational resources—particularly for large-scale datasets (>1 million cells). ChromFound bypasses these inefficiencies, enabling rapid integration of new data.
3. Scalability of genome-aware tokenization: The variable OCR tokenization ensures robust scalability across datasets with divergent OCR definitions. When integrating datasets with minimal OCR overlap (e.g., novel tissues or rare cell types), fixed peak sets discard biologically meaningful OCRs unique to new data. ChromFound dynamically encodes OCRs based on genomic coordinates, preserving dataset-specific regulatory elements critical for discovery-driven tasks.
Summary: ChromFound addresses two critical limitations of existing approaches: (1) dependency on raw fragment files and reprocessing workflows, and (2) inability to handle datasets with low OCR overlap. By genome-aware tokenization, our framework enables integration of heterogeneous scATAC-seq datasets without complex bioinformatic pipelines. This design aligns with the core objectives of foundation models—scalability, universality, and practical utility—as demonstrated by its state-of-the-art performance across diverse tasks.
Appendix F Analysis of Cell Type Annotation
F.1 Metric Value of Cell Type Annotation
To adapt ChromFound for cell type annotation, we leverage its four-layer pretrained model, extracting the encoder output as a general-purpose representation. The encoder’s hidden dimension is expanded from 128 to 256 and projected to a single dimension, forming a pooling tensor. This tensor is then processed by additional multilayer perceptron (MLP) layers to predict cell type logits. During fine-tuning, the pretrained decoder parameters are frozen, while gradients flow through the rest of the backbone to align with the classification objective. Dropout and LayerNorm are applied to mitigate overfitting and stabilize training, respectively. Benchmark methods are implemented using their default parameters as provided in their source codes.
For the experimental setup, we split the training data into 90% for training and 10% for validation. Training is conducted over 20 epochs using the AdamW optimizer with an initial learning rate of $5\times 10^{-4}$ and a 50-step learning rate warmup schedule. The model is selected based on validation performance and evaluated on the test set to compute accuracy and macro F1-score. For certain configurations, such as training on EPF_hydrop_2 and testing on VIB_10xv1_1, the learning rate is reduced to $2.5\times 10^{-4}$ to ensure stable convergence. All experiments are performed on a single machine with four NVIDIA A100 GPUs.
F.2 Confusion Matrix of Cell Type Annotation
In the cell classification task where EPF_hydrop_3 served as the training set and VIB_10xv1_2 as the test set, ChromFound achieves a 4.71% improvement in accuracy and a 14.69% increase in macro F1 score compared to the previous SOTA method, CellCano. To further elucidate the specific advancements of ChromFound, we present the confusion matrices of both methods for a more detailed comparative analysis.
Here are some biological insights from the comparison of confusion matrices:
(1) The observed misclassification rate of natural killer cells as cytotoxic T cells by the Cellcano model (19.35%) and its significant reduction by ChromFound (2.22%) highlights the importance of accurately distinguishing these functionally related yet distinct immune cell populations. Natural killer cells and cytotoxic T cells share cytotoxic properties, as both can mediate target cell lysis through perforin and granzyme pathways. However, they differ fundamentally in their ontogeny, activation mechanisms, and immune regulation.
(2) The high misclassification rate of CD16+ monocytes as CD14+ T cells by Cellcano (91.91%), and its substantial reduction by ChromFound (66%), underscores the challenge of accurately distinguishing myeloid from lymphoid lineages in cell type annotation based on scATAC-seq. CD16+ monocytes, a subset of non-classical monocytes, exhibit distinct chromatin accessibility patterns associated with Fc receptor signaling, inflammatory responses, and patrolling behavior, whereas CD14+ T cells, a less well-characterized subset, retain a predominantly lymphoid epigenetic signature. The improved classification by ChromFound suggests a refined ability to resolve lineage-specific regulatory elements, likely by better capturing differential enhancer accessibility and transcription factor binding landscapes unique to myeloid versus lymphoid cell fate.
(3) Cellcano misclassifies dendritic cells as CD14+ monocytes at a rate of 65.7% while ChromFound reduces this misclassification to 53.33%. The misclassification of dendritic cells as CD14+ monocytes is particularly interesting because both cell types share similar transcriptional signatures, especially in immune responses. dendritic cells and monocytes both play critical roles in antigen presentation and inflammation, which may lead to similarities in the chromatin accessibility profiles captured by scATAC-seq. From a biological perspective, this reduction in misclassification also underscores the importance of distinguishing between functionally distinct but phenotypically similar immune cell subsets. Dendritic cells, being key mediators of immune tolerance and initiation, have distinct regulatory networks compared to monocytes, which are more directly involved in inflammatory responses. By refining the accuracy of cell type annotation, ChromFound enables a more precise understanding of immune cell dynamics, particularly in the context of immune responses and disease progression.
Figure 5: Confusion matrix for cell type annotation task with a comparison of Cellcano and ChromFound.
Appendix G Supplementary results of denoising batch effect
We show the plots of the denoising batch effect as below. The results on To326K [68] from the Bone tissue are plotted in Fig. 6. The results on Kuppel139K [36] from the Heart tissue are plotted in Fig. 7. The results on Morabitol130K [49] from the Cortex tissue are plotted in Fig. 8. The results on PBMC169K [15] from the PBMC tissue are plotted in Fig. 9. In addition, the results of all metrics are listed in Table G.
Table 6: Comparison of Batch Correction and Biological Conservation Metrics Across Datasets. The best performance for each metric within each dataset is highlighted in bold.
Dataset Method Batch Correction Metrics Biological Conservation Metrics
$\text{ASW}_{\text{batch}}$ GraphConn $\text{ASW}_{\text{cell}}$ $\text{ARI}_{\text{cell}}$ $\text{NMI}_{\text{cell}}$
Cortex Morabito130K[49] batch 1/2 scVI[41] 0.8863 0.8595 0.5199 0.4537 0.4631
Scanorama[31] 0.9391 0.9211 0.5318 0.8234 0.7585
Harmony[35] 0.9174 0.9262 0.5793 0.8032 0.7892
scANVI[77] 0.8790 0.9697 0.6585 0.6185 0.7916
Liger[40] 0.8605 0.9598 0.5689 0.7899 0.7870
SCALEX[75] 0.9143 0.9274 0.5791 0.7896 0.7869
scBasset[85] 0.8052 0.9784 0.6361 0.6876 0.7230
ChromFound 0.9323 0.9808 0.6408 0.7531 0.8380
Bone To326K[68] batch 43/44 scVI[41] 0.9073 0.9283 0.4891 0.0070 0.0374
Scanorama[31] 0.8836 0.9580 0.5115 0.2325 0.3513
Harmony[35] 0.8508 0.9544 0.5539 0.2535 0.3066
scANVI[77] 0.7939 0.9506 0.5721 0.2835 0.4604
Liger[40] 0.8244 0.8874 0.5535 0.2167 0.3668
SCALEX[75] 0.7452 0.9506 0.5616 0.2835 0.4604
scBasset[85] 0.5163 0.7781 0.4339 0.0603 0.1035
ChromFound 0.8694 0.9885 0.6027 0.6426 0.6772
Heart Kuppe139K[36] donar av3/av10 scVI[41] 0.8128 0.8704 0.6247 0.8186 0.8355
Scanorama[31] 0.9378 0.7804 0.4707 0.6869 0.7072
Harmony[35] 0.8490 0.8407 0.5546 0.8066 0.8298
scANVI[77] 0.7844 0.7715 0.4337 0.5070 0.6553
Liger[40] 0.8374 0.8407 0.4196 0.4875 0.6287
SCALEX[75] 0.8638 0.8164 0.5424 0.7296 0.7711
scBasset[85] 0.7756 0.8282 0.6387 0.8055 0.7417
ChromFound 0.8589 0.8770 0.6155 0.9406 0.8980
PBMC 169K[15] HAR ddseq 1/2 scVI[41] 0.7452 0.8557 0.5794 0.5286 0.6888
Scanorama[31] 0.7964 0.8066 0.5613 0.5456 0.7005
Harmony[35] 0.7745 0.8120 0.5711 0.5525 0.7069
scANVI[77] 0.7480 0.8448 0.6223 0.5114 0.6773
Liger[40] 0.7755 0.8091 0.6197 0.5481 0.7020
SCALEX[75] 0.8221 0.7850 0.5534 0.5549 0.7003
scBasset[85] 0.7753 0.8266 0.5695 0.5456 0.7005
ChromFound 0.7923 0.8511 0.6443 0.5730 0.7156
Figure 6: The plots illustrate ChromFound’s superior performance in denoising batch effect on Bone To326K[68].
Figure 7: The plots illustrate ChromFound’s superior performance in denoising batch effect on Heart Kuppel139K [36].
Figure 8: The plots illustrate ChromFound’s superior performance in denoising batch effect on Cortex Morabitol130K [49].
Figure 9: The plots illustrate ChromFound’s superior performance in denoising batch effect on PBMC 169K[15].
Appendix H Benchmark Methods for Downstream Tasks
This section describes the benchmark methods used for evaluating the performance of ChromFound across various downstream tasks. These methods are selected based on their established effectiveness in single-cell data analysis, spanning cell clustering, cell type annotation, and cross-omics modality prediction.
H.1 Cell Clustering
Cell clustering methods aim to group single-cell data into meaningful clusters, often leveraging deep learning and generative models to handle the high dimensionality and sparsity of scATAC-seq data.
H.1.1 SCALE
SCALE [76] is a deep generative framework that employs probabilistic Gaussian Mixture Models to analyze high-dimensional scATAC-seq data. It outperforms existing tools in clustering by effectively capturing latent structures in the data.
H.1.2 SCALEX
SCALEX [75] is a deep learning method designed for integrating single-cell data. It projects cells into a batch-invariant embedding space in an online manner, enabling robust cell representation and achieving excellent clustering performance.
H.1.3 CASTLE
CASTLE [11] is a deep generative model tailored for single-cell epigenomic data. It uses a vector-quantized variational autoencoder framework [69] to extract discrete latent embeddings, excelling in cell clustering tasks by addressing the challenges of high dimensionality and sparsity.
H.1.4 scBasset
scBasset [85] is a sequence-based convolutional neural network method designed for modeling single-cell ATAC-seq data. By utilizing DNA sequence information from accessibility peaks and the expressive power of neural networks, scBasset achieves superior performance in tasks such as cell type identification and data integration across single-cell ATAC-seq and multiome datasets.
H.2 Denoising Batch Effect
Batch effect removal methods aim to align single-cell data across different batches while preserving biological signals, addressing technical variations that can obscure true biological differences. In this evaluation, we also benchmark SCALEX [75], previously introduced for cell clustering, and scBasset [85], noted for its sequence-based modeling, alongside other methods. These are included in the comparison for batch effect correction within their respective frameworks, with their detailed definitions provided in prior sections.
H.2.1 Harmony
Harmony [35] is a prominent batch correction technique in single-cell analysis, designed to harmonize datasets from diverse experimental batches. It employs an iterative strategy to align cells into a shared low-dimensional embedding, optimizing the similarity of cell types across batches while retaining biological variability, making it highly effective for downstream analyses in single-cell genomics.
H.2.2 scVI
scVI [41] is a package for end-to-end analysis of single-cell omics data. The package is composed of several deep generative models for omics data analysis.
H.2.3 Scanorama
Scanorama [31] is designed to be used in scRNA-seq pipelines downstream of noise-reduction methods, including those for imputation and highly-variable gene filtering. The results of Scanorama integration and batch correction can then be used as input to other tools for clustering, visualization, and analysis of scRNA sequences.
H.2.4 scANVI
scANVI[77] (single-cell ANnotation using Variational Inference; Python class SCANVI) is a semi-supervised model for single-cell transcriptomics data. In a sense, it can be seen as a scVI[41] extension that can leverage the cell type knowledge for a subset of the cells present in the data sets to infer the states of the rest of the cells. For this reason, scANVI can help annotate a data set of unlabelled cells from manually annotated atlases.
H.2.5 Liger
Liger[40] (Linked Inference of Genomic Experimental Relationships) integrates multi-omics single-cell data via integrative non-negative matrix factorization (iNMF) to define shared cell identities across protocols/species.
H.3 Cell Type Annotation
Cell type annotation methods focus on supervised learning approaches to assign cells to predefined classes, leveraging advanced neural network architectures for improved accuracy.
H.3.1 Cellcano
Cellcano [45] utilizes gene expression scores as input and employs a dual-phase training framework to achieve enhanced accuracy in cell type annotation across diverse datasets.
H.3.2 EpiAnno
EpiAnno [6] implements a probabilistic generative model with a Bayesian neural network. It delivers remarkable performance in cell type annotation, particularly when applied to diverse single-cell datasets.
H.4 Cross-omics Prediction
Cross-omics modality prediction methods aim to infer one modality (e.g., ATAC-seq) from another (e.g., RNA-seq), often using integrative neural network frameworks to model the relationships between modalities.
H.4.1 BABEL
BABEL [74] leverages an interoperable neural network model to translate between the transcriptome and chromatin profiles of individual cells, enabling effective cross-omics prediction.
H.4.2 CMAE
CMAE [82] learns a probabilistic coupling between different data modalities using autoencoders. It provides a robust framework for integrating and translating between single-cell data modalities.
H.4.3 scMoGNN
scMoGNN [72], an official winner in the overall ranking of modality prediction from the NeurIPS 2021 Competition, presents a general Graph Neural Network framework to facilitate multimodal single-cell data analysis, demonstrating superior performance in cross-omics prediction tasks.
H.4.4 scMM
scMM [48] is a novel deep generative model-based framework for single-cell multi-omics data analysis (e.g., transcriptome and chromatin accessibility). It employs a mixture-of-experts multimodal approach to extract interpretable joint representations and enable cross-modal generation, achieving end-to-end learning by modeling raw count data with modality-specific probability distributions.
H.5 Biological Application: Predicting Enhancer-Gene Link and Perturbation Response
This section outlines the benchmark methods evaluated for the biological application of predicting enhancer-gene links and perturbation responses using ChromFound. This task leverages single-cell data to infer regulatory relationships and responses to perturbations, with a focus on maintaining biological relevance. BABEL [74] and CMAE [82], the second best models in cross-omics prediction tasks, are included in this application aim to model the regulatory interactions between enhancers and genes,
Appendix I Evaluation Metrics for Downstream Tasks
This appendix details the evaluation metrics used for the downstream tasks in our study, organized by task. Each metric is defined mathematically to ensure clarity and reproducibility, with ranges and performance interpretations provided where applicable.
I.1 Cell Clustering
Cell clustering is an unsupervised learning task that groups single-cell data into clusters. When ground-truth labels (e.g., cell types) are available, we evaluate clustering performance using the following metrics.
I.1.1 Adjusted Rand Index (ARI)
The Rand Index (RI) measures the proportion of correctly grouped sample pairs:
$\text{RI}=\frac{\text{TP}+\text{TN}}{\text{TP}+\text{TN}+\text{FP}+\text{FN}},$
where TP, TN, FP, and FN denote the numbers of true positives, true negatives, false positives, and false negatives, respectively, based on whether pairs are correctly grouped together or apart in both the clustering and ground-truth labels. The Adjusted Rand Index (ARI) corrects for chance by adjusting the expected RI under random labeling:
$\text{ARI}=\frac{\text{RI}-\mathbb{E}[\text{RI}]}{\max(\text{RI})-\mathbb{E}[\text{RI}]},$
ranging from $-1$ to $1$, where higher values indicate better agreement between clustering results and ground-truth labels.
I.1.2 Fowlkes-Mallows Index (FMI)
The Fowlkes-Mallows Index (FMI) combines precision and recall for clustering:
$\text{FMI}=\sqrt{\frac{\text{TP}}{\text{TP}+\text{FP}}\cdot\frac{\text{TP}}{\text{TP}+\text{FN}}},$
where TP, FP, and FN are defined as above. It ranges from 00 to $1$, with higher values indicating better clustering performance.
I.1.3 Normalized Mutual Information (NMI)
NMI quantifies the shared information between clustering $C$ and ground-truth $Y$:
$\text{NMI}(Y,C)=\frac{2\cdot I(Y;C)}{H(Y)+H(C)},$
where $I(\cdot;\cdot)$ is mutual information and $H(\cdot)$ is entropy. It ranges from 00 to $1$, with higher values indicating greater alignment between the clustering and ground-truth labels.
I.1.4 Adjusted Mutual Information (AMI)
AMI adjusts NMI for chance by accounting for expected mutual information under random labeling:
$\text{AMI}=\frac{I(Y;C)-\mathbb{E}[I(Y;C)]}{\max(I(Y;C))-\mathbb{E}[I(Y;C)]},$
ranging from 00 to $1$, where higher values indicate a more robust agreement between clustering and ground-truth, especially in datasets with many clusters.
I.2 Denoising Batch Effect
Denoising Batch Effect aligns single-cell data across batches while preserving biological signals. We evaluate performance using biological conservation and batch mixing metrics implemented in [44].
I.2.1 Biological Conservation Metrics
We use the following metrics to assess the preservation of biological signals, where $C$ denotes the set of cell types.
Normalized Mutual Information (NMIcell)
$\text{NMI}_{\text{cell}}(Y,C)=\frac{2\cdot I(Y;C)}{H(Y)+H(C)},$
where $I(\cdot;\cdot)$ is mutual information and $H(\cdot)$ is entropy. It ranges from 00 to $1$, with higher values indicating better alignment with ground-truth cell types.
Adjusted Rand Index (ARIcell)
$\text{ARI}_{\text{cell}}=\frac{\text{RI}-\mathbb{E}[\text{RI}]}{\max(\text{RI})-\mathbb{E}[\text{RI}]},$
where $\text{RI}=\frac{\text{TP}+\text{TN}}{\text{TP}+\text{TN}+\text{FP}+\text{FN}}$ and TP, TN, FP, FN are defined as in the clustering section. It ranges from $-1$ to $1$, with higher values indicating better clustering agreement with cell types.
Average Silhouette Width (ASWcell)
$\text{ASW}_{\text{cell}}=\frac{1}{N}\sum_{i=1}^{N}\frac{b(i)-a(i)}{\max(a(i),b(i))},$
where $a(i)$ is the average distance of cell $i$ to others in its cell type, and $b(i)$ is the smallest average distance to another cell type. It ranges from $-1$ to $1$, with higher values indicating better preservation of biological clustering.
Average Biological Score (AvgBIO)
$\text{AvgBIO}=\frac{\text{ARI}_{\text{cell}}+\text{NMI}_{\text{cell}}+\text{ASW}_{\text{cell}}}{3},$
averaging the biological conservation metrics, ranging from 00 to $1$, with higher values indicating better preservation of biological signals.
I.2.2 Batch Mixing Metrics
We assess batch mixing using the following metrics, where $B$ denotes the set of batches.
Inverse Average Silhouette Width (ASWbatch)
$\text{ASW}_{\text{batch}}=1-\left|\frac{1}{N}\sum_{i=1}^{N}\frac{b(i)-a(i)}{\max(a(i),b(i))}\right|,$
where $a(i)$ and $b(i)$ are computed with respect to batch labels, with $a(i)$ as the average distance to other cells in the same batch and $b(i)$ as the smallest average distance to cells in another batch. It ranges from 00 to $1$, with higher values indicating better batch mixing.
Graph Connectivity (GraphConn)
$\text{GraphConn}=\frac{1}{|C|}\sum_{c\in C}\frac{|\text{LCC}(G_{c}^{\text{kNN}})|}{N_{c}},$
where $\text{LCC}(G_{c}^{\text{kNN}})$ is the size of the largest connected component in the k-nearest neighbors (kNN) graph of cells in cell type $c$, and $N_{c}$ is the number of cells in $c$. It ranges from 00 to $1$, with higher values indicating better connectivity across batches.
Average Batch Score (AvgBATCH)
$\text{AvgBATCH}=\frac{\text{ASW}_{\text{batch}}+\text{GraphConn}}{2},$
averaging the batch mixing metrics, ranging from 00 to $1$, with higher values indicating better batch correction.
I.3 Cell Type Annotation
Cell type annotation is a supervised task that assigns cells to ground-truth cell type. We use the following metrics to evaluate performance.
I.3.1 Accuracy
Accuracy measures the proportion of correctly classified cells:
$\text{Accuracy}=\frac{\sum_{c\in C}\text{tp}_{c}}{\sum_{c\in C}N_{c}},$
where $\text{tp}_{c}$ is the number of true positives for cell type $c$, and $N_{c}$ is the total number of cells in $c$. It ranges from 00 to $1$, with higher values indicating better classification performance.
I.3.2 Macro F1 Score
The Macro F1 Score averages the F1 scores across all cell types:
[…]
where $\text{Precision}_{c}=\frac{\text{tp}_{c}}{\text{tp}_{c}+\text{fp}_{c}}$ and $\text{Recall}_{c}=\frac{\text{tp}_{c}}{\text{tp}_{c}+\text{fn}_{c}}$, with $\text{fp}_{c}$ and $\text{fn}_{c}$ as false positives and false negatives for cell type $c$. It ranges from 00 to $1$, with higher values indicating better balanced performance across classes.
I.4 Cross-Omics Prediction
Cross-omics prediction infers one modality (e.g., ATAC-seq) from another (e.g., RNA-seq). We evaluate performance using correlation-based metrics.
I.4.1 Pearson Correlation Coefficient (PCC)
PCC measures the linear correlation between predicted ($x_{i}$) and true ($y_{i}$) values:
$\text{PCC}=\frac{\sum_{i=1}^{N}(x_{i}-\bar{x})(y_{i}-\bar{y})}{\sqrt{\sum_{i=1}^{N}(x_{i}-\bar{x})^{2}}\cdot\sqrt{\sum_{i=1}^{N}(y_{i}-\bar{y})^{2}}},$
where $\bar{x}$ and $\bar{y}$ are the means of the predicted and true values, respectively. It ranges from $-1$ to $1$, with higher (closer to $1$) values indicating stronger positive linear correlation.
I.4.2 Concordance Correlation Coefficient (CCC)
CCC extends PCC by accounting for bias and scale differences:
$\text{CCC}=\frac{2\cdot\rho\cdot\sigma_{x}\cdot\sigma_{y}}{\sigma_{x}^{2}+\sigma_{y}^{2}+(\mu_{x}-\mu_{y})^{2}},$
where $\rho$ is the PCC, $\sigma_{x}$ and $\sigma_{y}$ are the standard deviations of the predicted and true values, and $\mu_{x}$ and $\mu_{y}$ are their means. It ranges from $-1$ to $1$, with higher (closer to $1$) values indicating better agreement in both correlation and scale.
I.5 Predicting Enhancer-Gene Links and Perturbation Response
These tasks involve predicting regulatory relationships or perturbation outcomes, often treated as classification problems. We use the following metrics.
I.5.1 Area Under the Receiver Operating Characteristic Curve (AUC-ROC)
The Receiver Operating Characteristic (ROC) curve plots the True Positive Rate (TPR) against the False Positive Rate (FPR) at various classification thresholds:
$\text{TPR}=\frac{\text{TP}}{\text{TP}+\text{FN}},\quad\text{FPR}=\frac{\text{FP}}{\text{FP}+\text{TN}},$
where TP, TN, FP, and FN are true positives, true negatives, false positives, and false negatives, respectively. The Area Under the ROC Curve (AUC-ROC) quantifies the overall performance across all thresholds, ranging from 00 to $1$, with higher values indicating better classification performance, where $1$ represents perfect classification and $0.5$ represents random guessing.